OpenStudy (anonymous):

PLEASE HELP 1.) x+y+z=-1 2x-y+2z=-5 -x+2y-z=4

OpenStudy (anonymous):

I have -x-y-z=-1 2x+3y=3 those are my two separate answers now what?

hero (hero):

Just hang on a minute. I will solve this first, then explain.

OpenStudy (anonymous):

thank you!

hero (hero):

Okay, so here's what we can do here:

hero (hero):

1) Combine the 1st and 3rd equations together

hero (hero):

When you do that you will end up with 3y = 3 then y = 1

hero (hero):

Next, go ahead and replace y with 1 in each equation: x+1+z=-1 2x-1+2z=-5 -x+2(1)-z=4

hero (hero):

Notice that when you do that you'll end up with x + z = -2 x + z = -2 x + z = -2

OpenStudy (anonymous):

no u don't, I combined the 1st and 3rd equation and I got 2x+3y=3 ????

hero (hero):

Try to figure out how I got that.

hero (hero):

Each equation is the same

hero (hero):

I'll show you how to combine the 1st and 3rd equations

hero (hero):

x+y+z=-1 2x-y+2z=-5 -x+2y-z=4